OpenStudy (anonymous):

The multiplication of 3 digit number XY5 with digit Z yields X215. what is X+Y+Z???

OpenStudy (mathmath333):

by trial and error ,i found. \(\large\tt \begin{align} \color{black}{135\times 9=1215\\ x+y+z=1+3+9=13}\end{align}\)

OpenStudy (anonymous):

awesome. i thought there will be some other method like some sequential kind logical one :) thanks a ton @mathmath333

OpenStudy (mathmath333):

there can be equation method but that will be cumbersome ,notice here that z can only be odd digit except 1

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{z(100x+50y+5)=1000x+215}\end{align}\)

OpenStudy (anonymous):

ah exactly i was also thinking the same but couldn't get that far. thanks again :)

OpenStudy (anonymous):

true solving this equation is really messy :P

OpenStudy (mathmath333):

u can notice that their can be possible values of 215 can be when z=3 and y=0, and z=9 and y=3

OpenStudy (anonymous):

true for getting 5 at last we shld have z either 3 or 5 or 9, right...

OpenStudy (anonymous):

and for 215 we need the latter one. i get it